# BrightDreamer: Generic 3D Gaussian Generative Framework for Fast Text-to-3D Synthesis

### v1

@@ Abstract
Text-to-3D synthesis has recently seen intriguing advances by combining the text-to-image models with 3D representation methods, e.g., Gaussian Splatting (GS), via Score Distillation Sampling (SDS). However, a hurdle of existing methods is the low efficiency, per-prompt optimization for a single 3D object. Therefore, it is imperative for a paradigm shift from per-prompt optimization to one-stage generation for any unseen text prompts, which yet remains challenging. A hurdle is how to directly generate a set of millions of 3D Gaussians to represent a 3D object. This paper presents BrightDreamer, an end-to-end single-stage approach that can achieve generalizable and fast (77 ms) text-to-3D generation. Our key idea is to formulate the generation process as estimating the 3D deformation from an anchor shape with predefined positions. For this, we first propose a Text-guided Shape Deformation (TSD) network to predict the deformed shape and its new positions, used as the centers (one attribute) of 3D Gaussians. To estimate the other four attributes (i.e., scaling, rotation, opacity, and SH coefficient), we then design a novel Text-guided Triplane Generator (TTG) to generate a triplane representation for a 3D object. The center of each Gaussian enables us to transform the triplane feature into the four attributes. The generated 3D Gaussians can be finally rendered at 705 frames per second. Extensive experiments demonstrate the superiority of our method over existing methods. Also, BrightDreamer possesses a strong semantic understanding capability even for complex text prompts. The project code is available at https://vlislab22.github.io/BrightDreamer/.
Figure 1: A comparison between per-prompt optimization-based methods [52, 30, 10], and our single-stage generation-based approach with an end-to-end objective. (a) Optimization-based methods directly initialize a 3D representation model, e.g. 3D Gaussian Splatting (GS)  [22]. This process usually suffers from slow per-sample optimization (e.g., several hours for a single text). (b) By contrast, once trained, our approach directly generates 3D content for any unseen text prompt in 77 ms with a single run of a feedforward of our generator.
@@ 1 Introduction
Text-to-3D generation has recently received considerable attention in the computer graphics and vision community owing to its immersive potential across diverse applications, such as virtual reality and gaming [25].
@@
Recently, with the emergence of diffusion models [15, 47] and neural rendering techniques [37, 22], text-to-3D has witnessed an unprecedented technical advancement. In particular, pioneering methods, such as DreamFusion [43], LatentNeRF [36], SJC [55], have sparked significant interest in the research community, catalyzing a trend toward developing techniques for creating 3D assets from texts. The follow-up methods then focus on either quality improvement [46, 49, 56, 30] or geometry refinement [8, 31] or training efficiency [52, 64].
@@
The dominant paradigm of these methods is to randomly initialize a 3D representation model, e.g., Neural Radiance Fields (NeRF) [37] or Gaussian Splatting [22], and optimize such a model to align with a specific text prompt, as depicted in Fig. 1 (a). Unfortunately, these methods suffer from two critical challenges. Firstly, as per-prompt optimization usually requires several tens of thousands of iterations, this inefficiency brings a considerable obstacle to broader applications. It is significantly different from the mainstream training paradigm in the field of 2D image generation [50, 47] or 3D-aware image generation [48, 6, 17, 41, 5]: a generative model is trained with a collection of text-image pairs or images, and the model can generate the desired content from any input at the inference stage. Secondly, as demonstrated in Fig. 2(a), existing methods often fail to accurately process the complex texts. For example, the mainstream methods all fail to generate 3D content that input prompt contains complex interaction between multiple entities. This limitation arises from the models being trained on a single text prompt, which results in a degraded capability in comprehensive semantic understanding.
@@
Therefore, it is urgently needed for a paradigm shift from per-prompt optimization to develop a generic text-to-3D generation framework. Once trained, the generative framework should be able to generate content from any text prompts in the inference stage, as depicted in Fig. 1(b). Previously, some research efforts, e.g., ATT3D [33] and Instant3D [27], have been made grounded in NeRF representation. The core insight is to add a large number of texts and take them as conditional inputs to generate explicit spatial representations, such as triplane [5]. Nonetheless, in stark contrast to the volume rendering in NeRF, 3D GS representation for an object usually consists of millions of 3D Gaussians. Consequently, there exists an inherent and natural difficulty in converting the generation representations into 3D GS ones in their framework.
@@
Figure 2: DreamGaussian [52] and LucidDreamer [30] are both optimized for a single text. Our result is the direct generation. And for the display of our generalization, all the prompts do not appear in our training set. (a) is for showing the complex text understanding. (b) is for demonstrating our capability of generalization. It is noteworthy that light purple, deep purple, and light yellow don’t appear in the training set. (c) Interpolation between two prompts from color and shape perspective.
@@
In this paper, we propose BrightDreamer, an end-to-end single-stage framework that, for the first time, can achieve generalizable and fast (77 ms) text-to-3D GS generation. BrightDreamer exhibits a robust ability for complex semantic understanding (Fig. 2 (a)), and it demonstrates a substantial capacity for generalization (Fig. 2 (b)). In addition, same as traditional generative models [13], our generator can interpolate between two inputs (Fig. 2 (c)), which enables users to fully engage their imagination and creativity, expanding the potential for novel and nuanced design exploration. As stated before, the 3D GS representation of an object usually consists of several millions of 3D Gaussians. Consequently, directly generating such an extensive collection, is impractical. Our key idea is to redefine this generation problem as its equal problem, i.e., 3D shape deformation. Specifically, we place and fix some anchor positions to form the initial shape. Then, it can be deformed to the desired shape by giving different input prompts through our designed Text-guided Shape Deformation (TSD) network (Sec. 3.1). After the deformation, the new positions can be set to the centers of the 3D Gaussian. Upon establishing the basic shape, we elaborately design a Text-guided Triplane Generator (TTG) to generate a spatial representation of the 3D object (Sec. 3.2). Subsequently, we utilize the spatial feature of each center of 3D Gaussian to represent its whole feature and translate it into the remaining attributes (including scaling, rotation, opacity, and SH coefficient) through our well-designed Gaussian Decoder (Sec. 3.3). For TTG, grounded on our re-analysis of the previous convolution-based triplane generation process, we have identified and solved two primary deficiencies that necessitate rectification. One issue pertains to the spatial inhomogeneity observed during the calculation process, as illustrated in Fig. 4. The other issue arises from the single-vector style control mechanism similar to StyleGAN [19], which complicates the management of relationships between multiple entities.
@@ 2 Related Works
Text-to-3D Generation. Existing methods can be grouped into two categories. 1) Optimization-based methods typically commence with a randomly initialized 3D model, such as NeRF [37], and subsequently employ text-image priors [44, 47] to guide and optimize its parameters. After undergoing thousands of iterative refinements, this predefined 3D model progressively morphs to embody the shape described by the corresponding text input. DreamField [16] represents the inaugural foray into text-to-3D methodology, utilizing the pre-trained text-image model, CLIP [44], as a guiding mechanism for the optimization process of a predefined NeRF model. DreamFusion [43] proposes the Score Distillation Sampling (SDS) to transfer the prior of the 2D diffusion model [15] into a 3D representation model [37, 39], which achieves impressive performance and ignites the research enthusiastic for the text-to-3D task. VSD [56] and ISM [30] are devoted to re-designing the SDS loss [43], enabling much more local details of the 3D model. MVDream [49] and PerpNeg [2] attempt to solve the Janus problem, i.e., multi-face problem in some text prompts. 2) Generation-based methods, by contrast, aim to directly generate a 3D model from a given text, streamlining the process of text-to-3D generation. ATT3D [33] is the first attempt to train a NeRF model with multiple texts. Instant3D [27] designs some modules to map the text input to the EG3D model [5] and then use SDS to train this model. Nonetheless, different from the volume rendering in NeRF, 3D GS representation for an object usually consists of millions of 3D Gaussians. Therefore, it is inherently difficult to convert the NeRF representations into 3D GS ones. We propose BrightDreamer, a generic framework that, for the first time, can achieve fast (77 ms) text-to-3D GS generation. BrightDreamer exhibits a robust ability for complex semantic understanding (see Fig. 2).
@@
3D Gaussian Splatting (GS). Recently, 3D GS [22] has become the most popular 3D representation method of 3D objects or scenes. 3D GS shows a faster rendering speed and higher application potential than NeRF [37]. Within a short time, a large number of methods have been proposed to leverage 3D GS for diverse tasks, e.g., anti-aliasing novel view synthesis [65, 61], SLAM [60, 21, 35, 66], human reconstruction [26, 38, 24, 1, 28, 32], dynamic scene reconstruction [34, 63, 57, 62, 59], and 3D content generation [58, 10, 52, 64, 29, 30]. Our work is also based on 3D GS; however, we aim to develop a generic text-to-3D generation framework that can generate 3D Gaussians at a low latency (77 ms).
3D Generative Model. The 3D generative model can be divided into two categories. In a broad sense, NeRF [37], 3D GS [22], and their subsequent works [4, 39, 54, 7, 12, 65] can be seen as the novel view image generation from a set of input views. From a narrow perspective, the objective of 3D-aware image synthesis and 3D object generation is to directly generate the 3D representation itself, without any input views. For example, conditional NeRF [48, 6, 14] and conditional SDF [41, 17] take a latent code (Noise) as a condition to obtain different content. EG3D [5] and Instant3D [27] directly generate a triplane to represent the space. Differently, our BrightDreamer exclusively employs texts to train our 3D generative model, leveraging SDS. And it outputs a real 3D GS model, which is different from a conditional NeRF network. In addition, we introduce a better triplane design, which is different from EG3D and Instant3D.
@@
Figure 3: An overview of our BrightDreamer. The details of Spatial Transformer, ResConv Block and Upsample Block are shown in Fig. 5.
@@ 3 Method
Overview. The objective of BrightDreamer is to directly generate 3D Gaussians in response to text prompts. After training, BrightDreamer is capable of generating the 3D Gaussians with a remarkably low generation latency (about 77ms). And, the generated 3D Gaussians can be rendered at an impressive inference speed of over 700 frames per second. 3D Gaussians can be defined by five attributes, namely the center $p^{\prime}$, scaling $S$, rotation $R$, opacity $\alpha$, and SH coefficient $SH$. To directly generate 3D Gaussians, our key idea is two-fold: 1) Defining anchor positions, i.e., predefined positions, to estimate the center of 3D Gaussians; 2) Building implicit spatial representation, which can be decomposed to estimate the other four attributes of 3D Gaussians.
@@ 3.1 Text-guided Shape Deformation (TSD)
Anchor Position. The anchor position is predefined in this paper, which is a fixed coordinate. It serves as one of the inputs for the TSD. Specifically, we place the anchor positions on the vertices of a 3D grid, as represented by the gray points in Fig. 3. Then, we design the TSD network to predict their deviation to deform the initialized shape of the 3D grid, guided by the text prompt input.
@@
$score=softmax(\frac{W_{Q}(p)W_{K}(y)}{\sqrt{d}}),$ (1)
$h=score\cdot W_{V}(y),$ (2)
$output=FFN(h),$ (3)
where $p\in\mathbb{R}^{3}$ is the 3D coordinate of the anchor position, $y\in\mathbb{R}^{77\times 1024}$ is the text embedding of the input prompt, 77 and 1024 are the sentence length and the embedding dimension, $W_{Q}(\cdot)$, $W_{K}(\cdot)$ and $W_{V}(\cdot)$ are the query, key, value transformation function, $d$ is the feature dimension, $score$ represent the attention score between words and points, $h$ is the intermediate feature, $FFN(\cdot)$ is the feed-forward network.
The output of the TSD network is the offset $\Delta\in\mathbb{R}^{3}$ of the anchor positions. To ensure the stability of the training, we control the maximum extent to which a point can deviate from the anchor position. Specifically, given the degree of freedom $\beta\in\mathbb{R}$, we use the following equation to adjust the range of output into interval $(-\beta,\beta)$:
@@
Finally, we have the deformed position $p^{\prime}\in\mathbb{R}^{3}$, i.e. centers of 3D Gaussian, corresponding to input prompt, as follows:
@@ 3.2 Text-guided Triplane Generator (TTG)
Upon determining the centers of the 3D Gaussians, we need to obtain the other four attributes. To efficiently assign features to each Gaussian, the objective of TTG is to generate an implicit spatial representation in space, represented by the triplane. Therefore, we design a novel and highly efficient triplane generator, whose input is also the text prompts.
@@
Figure 4: The visualization of expanding 2D convolution kernel (cyan area) to 3D and its moving process in previous convolutional triplane generator [5]. We use $1\times 1$ convolutional kernel as an example. Only several positions are interacted, which leads to spatial inhomogeneous.
Another challenge is that given the complex prompts, squeezing a sentence into a single style feature vector to apply StyleGAN’s AdaIN operation [19, 20, 18] could result in a loss of local details. Therefore, we need a more fine-grained generation method guided by the word level, thus can retain more information of text encoder trained on large-scale dataset. Naturally, calculating cross-attention between the pixels of the feature map and words in the sentence is a better choice.
@@
ResConv Block. As illustrated in Fig. 5 (b), the Residual Convolutional Block consists of a sequence of components: a Layer Normalization layer, followed by the application of the SiLU activation function [11], and $3\times 3$ convolutional layers. Then a skip connection is established between input and output.
Upsample Block. As depicted in Fig. 5 (c), the Upsample Block begins with the interpolation of the feature map, enlarging it by a factor of 2×2$\times$2 ×. Following this upscaling, the enlarged feature map is further processed through a $3\times 3$ convolutional layer.
@@
It aims to obtain the other four attributes of 3D Gaussian for achieving the generation. Upon generating the triplane, comprised of the three feature planes $\pi_{xy}$, $\pi_{xz}$, $\pi_{yz}$, we can obtain the feature vector $\mathcal{F}\in\mathbb{R}^{32}$ of each Gaussian based on its center $p^{\prime}$. Subsequently, the feature vector needs to be converted into additional attributes of 3D Gaussian, including opacity $\alpha\in\mathbb{R}$, scaling $S\in\mathbb{R}^{3}$, rotation $R\in\mathbb{R}^{4}$, and SH coefficient.
Specifically, we first project the 3D coordinate onto the three planes, X-Y, X-Z, and Y-Z. Based on the projected 2D coordinates, we can derive the features $\mathcal{F}_{xy}$, $\mathcal{F}_{xz}$, $\mathcal{F}_{yz}$ according to the interpolation with their four vertex in the 2D feature maps. To ensure that the gradient back-propagation is evenly distributed across all three planes, we utilize an averaging operation to aggregate these features, thereby obtaining the 3D Gaussian’s feature $\mathcal{F}$. Given that the attributes of 3D Gaussian can be categorized into two groups, i.e., shape and color, we develop two distinct transformation modules $F_{shape}$ and $F_{color}$. Each module is a lightweight, two-layer Multi-Layer Perceptron (MLP) network. To enhance the gradient back-propagation to the TSD, the center of 3D Gaussian $p^{\prime}$ is additionally inputted into both modules:

### v2

@@ Abstract
Text-to-3D synthesis has recently seen intriguing advances by combining the text-to-image priors with 3D representation methods, e.g., 3D Gaussian Splatting (3D GS), via Score Distillation Sampling (SDS). However, a hurdle of existing methods is the low efficiency, per-prompt optimization for a single 3D object. Therefore, it is imperative for a paradigm shift from per-prompt optimization to feed-forward generation for any unseen text prompts, which yet remains challenging. An obstacle is how to directly generate a set of millions of 3D Gaussians to represent a 3D object. This paper presents BrightDreamer, an end-to-end feed-forward approach that can achieve generalizable and fast (77 ms) text-to-3D generation. Our key idea is to formulate the generation process as estimating the 3D deformation from an anchor shape with predefined positions. For this, we first propose a Text-guided Shape Deformation (TSD) network to predict the deformed shape and its new positions, used as the centers (one attribute) of 3D Gaussians. To estimate the other four attributes (i.e., scaling, rotation, opacity, and SH), we then design a novel Text-guided Triplane Generator (TTG) to generate a triplane representation for a 3D object. The center of each Gaussian enables us to transform the spatial feature into the four attributes. The generated 3D Gaussians can be finally rendered at 705 frames per second. Extensive experiments demonstrate the superiority of our method over existing methods. Also, BrightDreamer possesses a strong semantic understanding capability even for complex text prompts. The code is available in the project page.
Figure 1: A comparison between per-prompt optimization-based methods, and our feed-forward generation-based approach with an end-to-end objective. (a) Optimization-based methods directly initialize a 3D representation model, e.g. 3D Gaussian Splatting (GS). This process usually suffers from slow per-sample optimization (e.g., several hours for a single text). (b) By contrast, once trained, our approach directly generates 3D content for any unseen text prompt in 77 ms with a single run of a feed-forward of our generator.
@@ 1 Introduction
Text-to-3D generation has received considerable attention in the computer graphics and vision community owing to its immersive potential across diverse applications, such as virtual reality and video gaming [29]. Recently, with the emergence of diffusion models [17, 58] and neural rendering techniques [47, 26], text-to-3D has witnessed an unprecedented technical advancement. In particular, pioneering methods, such as DreamFusion [53], LatentNeRF [46], SJC [69], have sparked significant interest in the research community, catalyzing a trend toward developing techniques for creating 3D assets from texts. The follow-up methods then focus on either quality improvement [56, 62, 71, 37] or geometry refinement [8, 38] or training efficiency [65, 88].
@@
The dominant paradigm of these methods is to randomly initialize a 3D representation model, e.g., Neural Radiance Fields (NeRF) [47] or 3D Gaussian Splatting (3D GS) [26], and optimize such a model to align with a specific text prompt, as depicted in Fig. 1 (a). Unfortunately, these methods suffer from two critical constraints. Firstly, as per-prompt optimization usually requires several tens of thousands of iterations, this inefficiency brings a considerable obstacle to broader applications. It is significantly different from the mainstream training paradigm in the field of 2D image generation [63, 58, 57] or 3D-aware image generation [61, 5, 20, 51, 6]: a generative model is trained with a collection of text-image pairs or images, and the model can generate the desired content from any input at the inference stage. Secondly, as shown in Fig. 2(a), existing methods often fail to accurately process the complex texts. For example, the mainstream methods struggle in generating 3D content that input prompt contains complex interaction between multiple entities. This limitation arises from the models being trained on a single text prompt, resulting in a degraded capability in comprehensive semantic understanding.
@@
Therefore, it is urgently needed for a paradigm shift from per-prompt optimization to develop a generic text-to-3D generation framework. Once trained, the generative framework should be able to generate content from any text prompts in the inference stage, as depicted in Fig. 1(b). Furthermore, given the scarcity of 3D data in comparison to the abundance of 2D image data, leveraging well-trained 2D image diffusion models to facilitate the training of 3D generative models presents a more effective and resource-efficient approach. Previously, some research efforts, e.g., ATT3D [42] and Instant3D [33], have been made grounded in NeRF representation. The core insight is to add a large number of texts and take them as conditional inputs to generate explicit spatial representations, such as triplane [6]. Nonetheless, in stark contrast to the volume rendering in NeRF, 3D GS representation for an object usually consists of millions of 3D Gaussians. Consequently, there exists an inherent and natural difficulty in converting the generation representations into 3D GS ones in their framework.
@@
Figure 2: DreamGaussian [65] and LucidDreamer [37] are both optimized for a single text. Our result is the direct generation. And for the display of our generalization, all the prompts do not appear in our training set. (a) is for showing the complex text understanding. (b) is to demonstrate our capability of understanding details. It is noteworthy that light purple, deep purple, and light yellow don’t appear in the training set. (c) Interpolation between two prompts from color and shape perspectives.
@@
In this paper, we propose BrightDreamer, an end-to-end feed-forward framework that, for the first time, can achieve generalizable and fast (77 ms) text-to-3D GS generation. BrightDreamer exhibits a robust ability for complex semantic understanding (Fig. 2 (a)), and it demonstrates a substantial capacity for generalization (Fig. 2 (b)). Additionally, like traditional generative models [15], our generator can interpolate between two inputs (Fig. 2 (c)), allowing users to unleash their imagination and creativity, thus expanding the potential for novel and nuanced design exploration. As stated before, the 3D GS representation of an object usually comprises several hundreds of thousands of 3D GS elements. Thus, directly generating such a large collection is impractical. Our key idea is to redefine this generation problem as its equal problem, i.e., 3D shape deformation. Specifically, we place and fix some anchor positions to form the initial shape. Then, it can be deformed to the desired shape by giving different input prompts through our designed Text-guided Shape Deformation (TSD) network (Sec. 3.1). Then, the new positions can be set to the centers of the 3D Gaussian. Upon establishing the basic shape, we elaborately design a Text-guided Triplane Generator (TTG) to generate a spatial representation of the 3D object (Sec. 3.2). Subsequently, we utilize the spatial feature of each center of 3D Gaussian to represent its feature and translate it into the remaining attributes (including scaling, rotation, opacity, and SH coefficient) through our well-designed Gaussian Decoder (Sec. 3.3). For TTG, based on our re-analysis of the previous convolution-based triplane generation process, we identified and solved two primary deficiencies. One issue involves the spatial inhomogeneity observed during the calculation process, as shown in Fig. 4. The other issue arises from the single-vector style control mechanism similar to StyleGAN [22], which complicates managing relationships between multiple entities.
@@ 2 Related Works
Text-to-3D Generation. Existing methods can be grouped into two categories. 1) Optimization-based methods typically commence with a randomly initialized 3D model, such as NeRF [47], and subsequently employ text-image priors [54, 58] to guide and optimize its parameters. After undergoing thousands of iterative refinements, this predefined 3D model progressively morphs to embody the shape described by the corresponding text input. DreamField [19] represents the inaugural foray into text-to-3D methodology, utilizing the pre-trained text-image model, CLIP [54], as a guiding mechanism for the optimization process of a predefined NeRF model. DreamFusion [53] proposes the Score Distillation Sampling (SDS) to transfer the prior of the 2D diffusion model [17] into a 3D representation model [47, 50], which achieves impressive performance and ignites the research enthusiastic for the text-to-3D task. Inspired by this, numerous works [71, 70, 37, 67, 89, 36, 30, 84, 75] are devoted to re-designing the SDS loss, enabling much more local details of the 3D model. MVDream [62] and PerpNeg [2] attempt to solve the Janus problem, i.e., multi-face problem in some text prompts. Though these methods have great generalizability, it usually needs several hours to optimize a 3D model. 2) Generation-based methods, by contrast, aim to directly generate a 3D model from a given text, streamlining the process of text-to-3D generation. ATT3D [42] is the first attempt to train a NeRF model with multiple texts by SDS. Though its generalizability is limited, benefitting from SDS, the rendered image is better than previous methods [59, 7, 14, 41] and the generated content is not limited to 3D data [48, 60]. Instant3D [33] designs some modules to map the text input to the EG3D model [6] and then uses SDS to train this model. Concurrently, there are also some works focusing on text-to-image-to-3D model route [80, 66, 43, 73, 78, 72, 39, 85, 18, 93, 31, 16, 79]. However, due to the limited amount of 3D datasets, it is extremely hard to train a generic model to implement this target well (Sec. 4.3.4). Different from the mentioned works, we propose BrightDreamer, a generic framework that, for the first time, can achieve fast (77 ms) direct text-to-3DGS generation. Moreover, our method doesn’t depend on 3D data, which shows a greater potential.
@@
3D Gaussian Splatting (GS). Recently, 3D GS [26] has emerged as the leading method for 3D representation method of 3D objects or scenes, offering faster rendering speeds and greater application potential than NeRF [47]. In a short time, a large number of methods have been proposed to leverage 3D GS for diverse tasks, e.g., anti-aliasing novel view synthesis [90, 83], SLAM [82, 25, 45, 91], human reconstruction [34, 49, 28, 1, 32, 40], dynamic scene reconstruction [44, 86, 74, 87, 81], and 3D content generation [77, 10, 65, 88, 35, 37]. Our work builds on 3D GS, aiming to develop a generic text-to-3D generation framework that can generate 3D GS with low latency (77 ms).
@@
Figure 3: An overview of BrightDreamer. The details of Spatial Transformer, ResConv Block and Upsample Block are shown in Fig. 5.
@@ 3 Method
Overview. BrightDreamer aims to generate 3D Gaussians directly from text prompts. After training, it can generate the 3D GS model with a remarkably low generation latency (about 77ms). And, the generated 3D Gaussians can be rendered at an impressive inference speed of over 700 frames per second. Each 3D Gaussian is defined by five attributes: center $p^{\prime}$, scaling $S$, rotation $R$, opacity $\alpha$, and SH coefficient $SH$. Our key idea is two-fold: 1) Defining anchor positions, i.e., predefined positions, to estimate the center of 3D Gaussians; 2) Building implicit spatial representation, which can be decomposed to estimate the other four attributes of 3D Gaussians.
@@ 3.1 Text-guided Shape Deformation (TSD)
Anchor Position. The anchor positions are predefined positions, which are the fixed coordinates. It serves as one of the inputs for the TSD. Specifically, we place the anchor positions on the vertices of a 3D grid, as represented by the gray points in Fig. 3. Then, we design the TSD network to predict their deviation to deform the initialized shape of the 3D grid, guided by the text prompt input.
@@
$output=FFN(softmax(\frac{W_{Q}(p)W_{K}(y)}{\sqrt{d}})\cdot W_{V}(y)),$ (1)
where $p\in\mathbb{R}^{3}$ is the 3D coordinate of the anchor position, $y\in\mathbb{R}^{77\times 1024}$ is the text embedding of the input prompt, 77 and 1024 are the sentence length and the embedding dimension, $W_{Q}(\cdot)$, $W_{K}(\cdot)$ and $W_{V}(\cdot)$ are the query, key, value transformation function, $d$ is the feature dimension, $FFN(\cdot)$ is the feed-forward network. The output of the TSD network is the offset $\Delta\in\mathbb{R}^{3}$ of the anchor positions. To ensure the stability of the training, we control the maximum extent to which a point can deviate from the anchor position. Specifically, given the degree of freedom $\beta\in\mathbb{R}$, we use the following equation to adjust the range of output into interval $(-\beta,\beta)$:
@@
Finally, the deformed position $p^{\prime}\in\mathbb{R}^{3}$, which represents the centers of 3D Gaussian corresponding to input prompt, is formulated as follows:
@@ 3.2 Text-guided Triplane Generator (TTG)
After determining the centers of the 3D Gaussians, we need to obtain the other four attributes. To efficiently assign features to each Gaussian, the objective of TTG is to generate an implicit spatial representation in space, represented by the triplane. Thus, we design a novel and highly efficient triplane generator that takes text prompts as input.
@@
Figure 4: The visualization of expanding 2D convolution kernel (blue area) to 3D and its moving process in previous convolutional triplane generator [6]. We use $1\times 1$ convolutional kernel as an example. Only several positions are interacted, which leads to spatial inhomogeneous.
Another challenge is that given the complex prompts, squeezing a sentence into a single style feature vector to apply AdaIN [22, 23, 24] could result in a loss of local details. Therefore, we need a more fine-grained generation method guided by the word level, thus can retain more information of text encoder trained on large-scale dataset. Naturally, calculating cross-attention between the pixels of the feature map and words in the sentence is a better choice.
@@
ResConv Block. As shown in Fig. 5 (b), the Residual Convolutional Block includes a Layer Normalization layer, a SiLU activation function [12], and $3\times 3$ convolutional layers, with the skip connection between input and output.
Upsample Block. As depicted in Fig. 5 (c), the Upsample Block begins with interpolating the feature map to enlarge it by a factor of 2×2$\times$2 ×, followed by processing through a $3\times 3$ convolutional layer.
@@
We aim to obtain four additional attributes of 3D Gaussian necessary for generation. Once the triplane, consisting of the three feature planes $\pi_{xy}$, $\pi_{xz}$, $\pi_{yz}$, is generated, we can obtain the feature vector $\mathcal{F}\in\mathbb{R}^{32}$ of each Gaussian based on its center $p^{\prime}$. This feature vector is then converted into the additional attributes, including opacity $\alpha\in\mathbb{R}$, scaling $S\in\mathbb{R}^{3}$, rotation $R\in\mathbb{R}^{4}$, and SH coefficient.
Specifically, we first project the 3D coordinate onto three planes, X-Y, X-Z, and Y-Z. Based on the projected 2D coordinates, we can derive the features $\mathcal{F}_{xy}$, $\mathcal{F}_{xz}$, $\mathcal{F}_{yz}$ according to the interpolation with their four vertex in the 2D feature maps. To ensure the gradient back-propagation is evenly distributed across all three planes, we utilize an averaging operation to aggregate these features, thereby obtaining the 3D Gaussian’s feature $\mathcal{F}$. Given that the attributes of 3D Gaussian can be categorized into two groups, i.e., shape and color, we develop two distinct transformation modules $F_{shape}$ and $F_{color}$. Each module is a lightweight, two-layer Multi-Layer Perceptron (MLP) network. To enhance the gradient back-propagation to the TSD, the center of 3D Gaussian $p^{\prime}$ is additionally inputted into both modules: